### v2

@@ Abstract
Cyber-physical systems like autonomous vehicles are tested in simulation before deployment, using domain-specific programs for scenario specification. To aid the testing of autonomous vehicles in simulation, we design a natural language interface, using an instruction-following large language model, to assist a non-coding domain expert in synthesising the desired scenarios and vehicle behaviours. We show that using it to convert utterances to the symbolic program is feasible, despite the very small training dataset. Human experiments show that dialogue is critical to successful simulation generation, leading to a 4.5 times higher success rate than a generation without engaging in extended conversation.
@@ 1 Introduction
Testing autonomous vehicles exclusively on public roads, especially in near-crash scenarios, is not feasible [1]. Iterative development and testing in simulation is essential [2], especially in early stages of the development process. Simulators like CARLA [3] support this iterative development process for autonomous vehicles (AVs). Using CARLA, engineers can generate driving scenarios using Scenic [4]: probabilistic scenario description languages such as Scenic specify a distribution over the driving scenarios that satisfy the description, and then sample many simulation instances for evaluation. In this way, challenging driving scenarios can be generated [5, 6], which can then be used for evaluation of control algorithms for autonomous vehicles without a physical system, safely and cost-effectively.
@@
Writing programs to specify driving scenarios in Scenic, or more generally programming in domain-specific languages (DSLs), is challenging with steep learning curves. Domain experts with deep knowledge of edge cases and desired driving behaviours may lack proficiency in Scenic or other DSLs, or even when they do have such proficiency, they may still benefit from some assistance in their workflow. A natural language interface in the form of a dialogue system in which engineers converse with a chatbot that knows how to program in Scenic is a useful tool for facilitating this. In this project, we aim to develop such a dialogue system, and by doing so increase engineer access to simulation-driven testing, by allowing them to interactively synthesise scenarios using natural language dialogue.
@@
Figure 1: Dialogue System Overview. It facilitates driving scenario generation using natural language in code generation (§ 3.1) and conversation (§ 3.2) modes.
@@
This is a challenging task for two reasons. Firstly due to data scarcity (only 32 examples of Scenic programs with natural language descriptions are available online), approaches in developing a dialogue system using instruction-following large language models (IFLLM) [7] are not directly applicable. Secondly, the scenario description provided by the user may be underspecified: missing some details about the desired driving scenario in the description that is only exposed through interaction with the simulator. Figure 1 illustrates this phenomenon. When a user utters “an ego vehicle overtakes the two adversary vehicles”, the dialogue system can synthesize the simulation instance satisfying the description, but the description itself does not provide all information about the situation, which in this case is the manner of overtaking (could be one by one, or both at the same time). The user only becomes aware of the necessity to provide such details by observing simulation instances. In this context, it would be helpful for natural language interface to allow the user to express feedback to the system, which can then be used to update the program and in turn generate simulation instances.
@@
To this end, this project makes the following contributions: 1) we create a dataset of English description-Scenic program pairs for a variety of driving scenarios; 2) we investigate how IFLLM can be used to synthesize Scenic programs from natural language expressions in an embodied conversation where the user converses with a chatbot and reacts to driving simulation instances; and 3) we conduct human trials to evaluate the dialogue system, and in particular the value of having multiple dialogue turns. Our results show that the dialogue system can be a valuable tool for facilitating interactive driving scenario generation.
@@ 2 Related Work
The problem of converting natural language to executable programs has been widely studied as semantic parsing [8, 9, 10]. Recently, using IFLLMs has been explored for this task and also the wider problem of arbitrary code generation, including auto-completion [11] to assist the software development process. In this project, we concentrate on developing a dialogue system that generates Scenic programs using LLMs. It is a low-resource domain where only a handful of examples are available. This is in contrast to other popular programming languages like Python, which benefit from many code repositories for training [12]. Furthermore, the code for high-resource programming languages follows style guidelines, making it easier for models to learn generalisations from the well-structured data. In contrast, the Scenic program repository was written in different styles by different software engineers.
@@
Testing is a fundamental process within software engineering, which is particularly important for designing safety-critical systems, such as AVs [13, 14]. There are a variety of methods for testing and validating AV design, even with black-box components [15]. This project focuses on testing by generating driving scenarios, as a complementary procedure to formal verification. Such testing involves interaction with expert users, which is essential for human-centred design [16]. Driving scenario generation has been explored before: both direct scenario generation [17, 18] or by utilizing Scenic [19, 20]. What is unique about this project is: (i) human experiments; and (ii) utilizing dialogue so as to align the generated driving scenarios with the user’s intent. We view this latter contribution as essential, given the ubiquitous phenomenon of natural language pragmatics [21]: speakers of natural language often intend to convey a meaning that goes beyond what they make linguistically explicit, relying instead on their interlocutor’s capacity to decode the intended content using the linguistic and non-linguistic context. There is no guarantee that the sampling processes in Scenic match how competent interlocutors perform this task, and so errors in decoding the intended message via Scenic sampling may happen, prompting the need for user feedback.
@@
Natural language interfaces in the form of a dialogue system have become a popular means for users to interact with systems without them needing to know the underlying system nor how to program it[22, 23]. Broadly, dialogue systems are categorized by their function as goal-oriented or for chit-chat. We are designing a goal-oriented dialogue system with tool use [24], which in this case is the CARLA simulator. What is unique in our project is that the driving scenarios procured by the simulator are not observed in the dialogue system and are latent, with observations coming just from the user’s feedback after observing simulation instances.
@@ 3.1 Code Generation
To generate a program $\hat{p}$ from natural language description $d$, we constructed a dataset $\mathcal{D}=\{(d_{i},p_{i})\}_{i=1}^{N}$ of size $N=105$ description-program pairs, via manual augmentation of the few exemplars available on the web. [^1] https://github.com/BerkeleyLearnVerify/Scenic/tree/2.x/examples Additionally in our manual augmentation process we standardize the codebase to minimize variability and ease the learning process in this low-resource regime. Each program is standardised in its coding style, with five distinct code blocks—map and model, constants, agent’s behaviour, spatial relations, and scenario specification—to minimize the variation between the exemplars. The dataset $\mathcal{D}$ is indexed via embeddings of descriptions. [^2] https://huggingface.co/BAAI/bge-small-en-v1.5
@@
The description with the retrieved exemplars are then used to create the $\mathtt{CODE}$ prompt (see Figure 3 for prompt specification) which is then used to generate the program $\hat{p}$ using an $\mathtt{IFLLM}$:
@@
There is no guarantee that the program $\hat{p}$ will execute. To fix this, we perform error feeding: the program $\hat{p}$ is passed to the simulator to attempt to create simulation instances; and if an error occurs, the natural language description, $\hat{p}$ and its errors are passed back to the $\mathtt{IFLLM}$, to attempt to self-correct the output. If after three attempts of error feeding the generated program doesn’t execute, the user is asked to provide an alternative description.
@@ 3.2 Conversation
When $\hat{p}$ executes, $n=3$ instances of the driving scenario simulation are shown to the user, who is given the opportunity to react to these, either by indicating successful generation or by providing natural language feedback $f$ that corrects or refines the results, in an attempt to provide simulations that better align with the user’s communicative intent.
Feedback is a response to the simulations and not to $\hat{p}$ (which is not observable to the user): it can feature and refer to information that is not present in the Scenic program but is present in the simulation due to the nature of random sampling. Because of this, feeding only $f$ to $\mathtt{IFLLM}$ is suboptimal due to its missing context, which cannot be given in a principled way (e.g. conditioning on the simulation video or some domain-specific trace would lead to further data scarcity, which is already problematic in this domain). To alleviate this issue, we feedback an updated description by summarizing $d$ and $f$, using a $\mathtt{SUMM}$ prompt (see the Appendix for prompt specification), to generate the updated description $d_{1}$. With this description, our aim is to guide $\mathtt{IFLLM}$, minimize the overall context, and not introduce hallucinations and unwanted behaviours.
@@
The conversation continues up to $m=4$ turns. If the user is still not satisfied with the simulations, the conversation is evaluated as unsuccessful.
@@ 4.2 Human Evaluation
Figure 5: Commutative scenario generation success over the number of dialogue turns with (5(a)) and without (5(b)) execution errors. E.g, if on the 3rd turn, the system fails to produce executable code, and the user is asked to paraphrase leading to successful scenario generation, we count this as success on 4 dialogue turns with execution errors and 3 dialogue turns without execution errors. The results show a positive correlation between the successful scenario generation and the number of dialogue turns for all types of scenarios considered.
@@ 4.2.1 Experimental Setup
We perform a human evaluation of the dialogue system. Visual stimuli used for human trials is given in Figure 4. The user is presented with a (diagrammatic) stimulus and asked to describe it. After that, CARLA is run to produce 3 simulation instances. [^3] The dialogue system can produce many (different) instances, but to keep the stimuli contained, we limit it to 3. After observing it, the user may respond with further natural language input to what they see, or check “satisfied with the scenario produced” and move to the next stimulus. The user can have up to $m=4$ dialogue turns. If code generation does not produce executable code, the user is asked to paraphrase their description as if it is a fresh interaction. The stimuli are categorised into three types: bypassing (e.g. the user stimulus may yield the user description “an ego overtakes an adversary”), intersections (e.g. “an ego goes left at a 4-way intersection while the adversary approaches from behind.”), and pedestrians (e.g.“an ego goes left and yields to the pedestrian”). The experiments were conducted with 20 users who have a driving license (as a proxy for domain proficiency), with each asked to describe 5 of the 25 stimuli. In total, 100 conversations were recorded: 44 intersections, 19 bypassing, and 37 pedestrian scenarios.
@@ 4.2.2 Results and Discussion
Figure 5 outlines the results of human evaluation. As the number of dialogue turns increases so does the the percentage of successfully generated scenarios: without dialogue, only 6% of the code generations were successful, while conversations of up to four turns raise this to 28%. This holds per scenario type, with or without execution errors. This suggests that the extended conversation enables successful scenario generation. Table 2 gives examples of a successful and failed scenario generation. Error analysis shows that failed scenario generation conversations had one or more of these errors:
  • Execution error (62.5%): the code produced does not execute and the user is asked to paraphrase one or more times. This happens due to a missing correspondence between Scenic API and the description. E.g. the description features “T-junction”, while the Scenic API would expect 3-way intersection, and the code features junction as a primitive rather than intersection.
@@ 5 Conclusions
We have presented a dialogue system for conversational driving scenario generation. Our results show that users can in principle generate desired driving scenarios with conversation being a crucial element in reducing scenario generation errors.
For future work, we envision several developments. Firstly, there is a gap of failed to generate scenarios even with the extended conversation. To address this, further techniques could be considered, including data augmentation [35], constrained decoding [36], prompt engineering [37], and memory-based techniques for explicit state tracking [38]. Second, the current setup does not dynamically change the controller and use the interface to test it. This should be evaluated and considered testing of such new controllers in tandem to formal verification methods [39].

### v3

@@ Abstract
Cyber-physical systems like autonomous vehicles are tested in simulation before deployment, using domain-specific programs for scenario specification. To aid the testing of autonomous vehicles in simulation, we design a natural language interface, using an instruction-following large language model (LLM), to assist a non-coding domain expert in synthesising the desired scenarios and vehicle behaviours. We show that using it to convert utterances to the symbolic program is feasible, despite the very small training dataset. Human experiments show that dialogue is critical to successful simulation generation, leading to a 4.5 times higher success rate than generation without engaging in extended conversation.
@@ 1 Introduction
Testing autonomous vehicles exclusively on public roads, especially in near-crash scenarios, is not feasible [1]. Iterative development and testing in simulation is essential [2], especially in the early stages of the development process. Simulators like CARLA [3] support this iterative development process for autonomous vehicles (AVs). Using CARLA, engineers can generate driving scenarios using Scenic [4]: probabilistic scenario description languages such as Scenic specify a distribution over the driving scenarios that satisfy the description, and then sample many simulation instances for evaluation. In this way, challenging driving scenarios can be generated [5, 6], which can then be used for evaluation of control algorithms for autonomous vehicles without a physical system, safely and cost-effectively.
@@
Writing programs to specify driving scenarios in Scenic, or more generally programming in domain-specific languages (DSLs), is challenging with steep learning curves. Domain experts with deep knowledge of edge cases and desired driving behaviours may lack proficiency in Scenic or other DSLs, or even when they do have such proficiency, they may still benefit from some assistance in their workflow. A natural language interface in the form of a dialogue system in which engineers converse with a chatbot that knows how to program in Scenic is a useful tool for facilitating this. In this paper, we aim to develop such a dialogue system, and by doing so increase engineer access to simulation-driven testing, by allowing them to interactively synthesise scenarios using natural language dialogue.
@@
Figure 1: Dialogue System Overview. It facilitates driving scenario generation using natural language in code generation (§ 3.1) and conversation (§ 3.2).
@@
This is a challenging task for two reasons. Firstly due to data scarcity (only 32 examples of Scenic programs with natural language descriptions are available online), approaches in developing a dialogue system using instruction-following large language models (IFLLM) [7] are not directly applicable. Secondly, the scenario description provided by the user may be underspecified: missing some details about the desired driving scenario in the description that is only exposed through interaction with the simulator. Figure 1 illustrates this phenomenon. When a user utters “an ego vehicle overtakes the two adversary vehicles”, the dialogue system can synthesize the simulation instance satisfying the description, but the description itself does not provide all information about the situation, which in this case is the manner of overtaking (could be one by one, or both at the same time). The user only becomes aware of the necessity to provide such details by observing simulation instances. In this context, it would be helpful for a natural language interface to allow the user to express feedback to the system, which can then be used to update the program and, in turn, generate simulation instances.
@@
To this end, this paper makes the following contributions: 1) we create a dataset of English description-Scenic program pairs for a variety of driving scenarios; 2) we investigate how IFLLM can be used to synthesize Scenic programs from natural language expressions in an embodied conversation where the user converses with a chatbot and reacts to driving simulation instances; and 3) we conduct human trials to evaluate the dialogue system, and in particular the value of having multiple dialogue turns. Our results demonstrate that the dialogue system can be a valuable tool for facilitating the generation of interactive driving scenarios.
@@ 2 Related Work
The problem of converting natural language to executable programs has been widely studied as semantic parsing [8, 9, 10]. Recently, using IFLLMs has been explored for this task and also the wider problem of arbitrary code generation, including auto-completion [11] to assist the software development process. In this paper, we concentrate on developing a dialogue system that generates Scenic programs using LLMs. It is a low-resource domain where only a handful of examples are available. This is in contrast to other popular programming languages like Python, which benefit from many code repositories for training [12]. Furthermore, the code for high-resource programming languages follows style guidelines, making it easier for models to learn generalisations from the well-structured data. In contrast, the Scenic program repository was written in different styles by different software engineers.
@@
Testing is a fundamental process within software engineering, which is particularly important for designing safety-critical systems, such as AVs [13, 14]. There are a variety of methods for testing and validating AV design, even with black-box components [15]. This paper focuses on testing by generating driving scenarios as a complementary procedure to formal verification. Such testing involves interaction with expert users, which is essential for human-centred design [16]. Driving scenario generation has been explored before: both direct scenario generation [17, 18] or by utilising Scenic [19, 20]. What is unique about this work is: (i) human experiments; and (ii) utilising dialogue to align the generated driving scenarios with the user’s intent. We view this latter contribution as essential, given the ubiquitous phenomenon of natural language pragmatics [21]: speakers of natural language often intend to convey a meaning that goes beyond what they make linguistically explicit, relying instead on their interlocutor’s capacity to decode the intended content using the linguistic and non-linguistic context. There is no guarantee that the sampling processes in Scenic match how competent interlocutors perform this task, and so errors in decoding the intended message via Scenic sampling may happen, prompting the need for user feedback.
@@
Natural language interfaces in the form of a dialogue system have become a popular means for users to interact with systems without needing to know the underlying system or how to program it[22, 23]. Broadly, dialogue systems are categorised by their function as goal-oriented or for chit-chat. We are designing a goal-oriented dialogue system with tool use [24], which in this case is the CARLA simulator. What is unique about our work is that the driving scenarios generated by the simulator are not observed in the dialogue system and are latent, with observations coming only from the user’s feedback after observing simulation instances.
@@ 3.1 Code Generation
To generate a program $\hat{p}$ from natural language description $d$, we constructed a dataset $\mathcal{D}=\{(d_{i},p_{i})\}_{i=1}^{N}$ of size $N=105$ description-program pairs [^1] https://huggingface.co/datasets/assistive-autonomy/scenic-driving-scenarios, via manual augmentation of the few exemplars available on the web. [^2] https://github.com/BerkeleyLearnVerify/Scenic/tree/2.x/examples Additionally, in our manual augmentation process, we standardise the codebase to minimise variability and ease the learning process in this low-resource regime. Each program is standardised in its coding style, with five distinct code blocks—map and model, constants, agent’s behaviour, spatial relations, and scenario specification—to minimise the variation between the exemplars. The dataset $\mathcal{D}$ is indexed via embeddings of descriptions. [^3] https://huggingface.co/BAAI/bge-small-en-v1.5
@@
The description with the retrieved exemplars are then used to create the $\mathtt{CODE}$ prompt (see Figure 3 for the prompt specification) which is then used to generate the program $\hat{p}$ using an $\mathtt{IFLLM}$:
@@
There is no guarantee that the program $\hat{p}$ will execute. To fix this, we perform error feeding: the program $\hat{p}$ is passed to the simulator to attempt to create simulation instances; and if an error occurs, the natural language description, $\hat{p}$ and its errors are passed back to the $\mathtt{IFLLM}$, to attempt to self-correct the output. If, after three attempts of error feeding, the generated program doesn’t execute, the user is asked to provide an alternative description.
@@ 3.2 Conversation
When $\hat{p}$ executes, $n=3$ instances of the driving scenario simulations are shown to the user, who is given the opportunity to react to these, either by indicating successful generation or by providing natural language feedback $f$ that corrects or refines the results, in an attempt to provide simulations that better align with the user’s communicative intent.
Feedback is a response to the simulations and not to $\hat{p}$ (which is not observable to the user): it can feature and refer to information that is not present in the Scenic program but is present in the simulation due to the nature of random sampling. Because of this, feeding only $f$ to $\mathtt{IFLLM}$ is suboptimal due to its missing context, which cannot be given in a principled way (e.g. conditioning on the simulation video or some domain-specific trace would lead to further data scarcity, which is already problematic in this domain). To alleviate this issue, we feedback an updated description by summarising $d$ and $f$, using a $\mathtt{SUMM}$ prompt (see the Figure 3 for the prompt specification), to generate the updated description $d_{1}$. With this description, our aim is to guide $\mathtt{IFLLM}$, minimise the overall context, and not introduce hallucinations and unwanted behaviours.
@@
The conversation continues up to $m=4$ turns. If the user remains unsatisfied with the simulations, the conversation is deemed unsuccessful.
@@ 4.2 Human Evaluation
Figure 5: Commutative scenario generation success over the number of dialogue turns in human evaluation with (5(a)) and without (5(b)) execution errors. For example, if, on the 3rd turn, the system fails to produce executable code and the user is asked to paraphrase, leading to successful scenario generation, we count this as success on 4 dialogue turns with execution errors and 3 dialogue turns without execution errors. The results show a positive correlation between the successful scenario generation and the number of dialogue turns for all types of scenarios considered.
@@ 4.2.1 Experimental Setup
We perform a human evaluation of the dialogue system. Visual stimuli used for human trials are given in Figure 4. The user is presented with a (diagrammatic) stimulus and asked to describe it. After that, CARLA is run to produce 3 simulation instances. [^4] The dialogue system can produce many (different) instances, but to keep the stimuli contained, we limit it to 3. After observing it, the user may respond with further natural language input to what they see, or check “satisfied with the scenario produced” and move to the next stimulus. The user can have up to $m=4$ dialogue turns. If code generation does not produce executable code, the user is asked to paraphrase their description as if it is a fresh interaction. The stimuli are categorised into three types: bypassing (e.g. the user stimulus may yield the user description “an ego overtakes an adversary”), intersections (e.g. “an ego goes left at a 4-way intersection while the adversary approaches from behind.”), and pedestrians (e.g.“an ego goes left and yields to the pedestrian”). The experiments were conducted with 20 users who have a driving license (as a proxy for domain proficiency), with each asked to describe 5 of the 25 stimuli. In total, 100 conversations were recorded: 44 intersections, 19 bypassing, and 37 pedestrian scenarios.
@@ 4.2.2 Results and Discussion
Figure 5 outlines the results of human evaluation. As the number of dialogue turns increases, so does the percentage of successfully generated scenarios: without dialogue, only 6% of the code generations were successful, while conversations of up to four turns raise this to 28%. This holds per scenario type, with or without execution errors. This suggests that the extended conversation enables successful scenario generation. Table 2 gives examples of a successful and failed scenario generation. Error analysis shows that failed scenario generation conversations had one or more of these errors:
  • Execution error (62.5%): the code produced does not execute, and the user is asked to paraphrase one or more times. This happens due to a missing correspondence between Scenic API and the description. E.g. the description features “T-junction”, while the Scenic API would expect 3-way intersection, and the code features junction as a primitive rather than intersection.
@@ 5 Conclusions
We have presented a dialogue system for generating conversational driving scenarios. Our results show that users can, in principle, generate desired driving scenarios with conversation being a crucial element in reducing scenario generation errors.
For future work, we envision several developments. Firstly, there is a gap in generating scenarios, despite the extended conversation. To address this, further techniques could be considered, including data augmentation [35], constrained decoding [36], prompt engineering [37], and memory-based techniques for explicit state tracking [38]. Second, the current setup does not dynamically change the controller and use the interface to test it. This should be evaluated and considered testing of such new controllers in tandem with formal verification methods [39].